[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jul 27, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 27 20:25:43 UTC 2017


HSU Shift Log for Thu Jul 27 2017 05:47:55 GMT-0400 (EDT)

Shift: A

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None.

AWIPS-2/NCP issues: One CAVE session on the AWIPS-2 at the JHT desk refused  
to respond to the pan/zoom of the mouse. Another CAVE session on the same  
workstation had no such problem.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170727/9974ee8a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list